A ball is thrown vertically upward with an initial velocity of 3 m/s from the window of a tall building. The ball strikes the ground 4 seconds later. Determine the velocity at which the ball strikes the groun
1
Expert's answer
2020-05-25T11:16:57-0400
"v_y=v_0+gt"

"v_y=-3+(9.8)(4)=36\frac{m}{s}"

Direction: downwards.